OpenStudy (anonymous):

The sum of two numbers is 15. The difference of the first number and three will equal the second number. Write and solve the system of equations using the substitution method.

OpenStudy (homeworksucks):

>The sum of two numbers Lets call them x and y so x+y=15. >The difference of the first number and 3 equals the second number So x-3=y

OpenStudy (homeworksucks):

Your finished system is\[x+y=15\] \[x-3=y\]

OpenStudy (homeworksucks):

So in order to substitute, solve for one variable in one equation. Since in the second equation its already done for us, that makes life easier. So you get x+y=15, right and you know that y=x-3. So plug in x-3 for y So x+(x-3)=15 2x-3=15 From there you can solve for x 2x-3=15 2x-3+3=15+3 2x=18 2x/2=18/2 x=9

OpenStudy (homeworksucks):

But, we still have to solve for y. We know that x-3=y right? So just plug in 9 for x and chug away. x-3=y 9-3=y y=6

OpenStudy (homeworksucks):

So our solution in (x,y) form is (9,6). We can double check this by doing some more plugging in. We know that x+y=15, so 9+6 has to equal 15, which it does, so we know its right.

OpenStudy (homeworksucks):

I hope this helps

OpenStudy (anonymous):